OpenStudy (anonymous):

using de Moivre's theorem to find (1-i)^6 in the form a+bi (begin by expressing 1-i in pola forms)

OpenStudy (helder_edwin):

do u know how to find the polar form of a complex number?

OpenStudy (anonymous):

no, i havent learnt it yet

OpenStudy (helder_edwin):

if \(z=a+ib\), its polar form is \[\large z=\rho(\cos\theta+i\sin\theta) \] where \(\rho=|z|=\sqrt{a^2+b^2}\) and \(\theta\) is the angle between the positive x-axis and the vector \(z\).

OpenStudy (helder_edwin):

so \[\large \rho=\sqrt{1^2+(-1)^2}=\sqrt{2} \]

OpenStudy (helder_edwin):

the angle (argument) can be measured either counterclockwise or, clockwise if z is in the III or IV quadrants or counterclockwise if < is in the I or II quadrants. how were u taught to do this?

OpenStudy (helder_edwin):

* z not <